Skip to main content

AAVSOnet status, 2018-03-15

1 post / 0 new
HQA
HQA's picture
AAVSOnet status, 2018-03-15

There have been some recent and ongoing failures of which all users should be aware.

The HQ processing computer (occam) has had a software failure during an operating system update.  Problems started about March 2, and finally resulted in complete shutdown on March 10.  The cause of the problem is known, and possible solutions are being studied.  It is likely that the computer will be back on-line early next week, with new data being available to researchers a few days later after the raw images are transferred to HQ.  Note that the ftp site, ftp.aavsonet.aavso.org, is also hosted on occam, and so is currently unavailable.

Listed below are hardware or operational problems for individual systems.

BSM_S  is currently off-line for a few days.  Peter has installed the new filter wheel, and will be obtaining filter offsets for the new filters (Sloan griz plus diffraction grating) when he returns.  BSM_S was down earlier this month due to cracked bearings on the focuser, but Peter replaced those bearings on March 9.

BSM_NH is currently off-line.  A sudden warming melted snow and soaked some connectors that were on the floor of the enclosure.  While cleaning that issue up, it was noticed that the focus controller had failed, due to water that had entered its case sometime in the past and finally corroded the circuit board.  All told, several cables and a new controller had to be ordered.  With the ongoing snow in New Hampshire, finding dry weather to finish and test the repairs has been a challenge.  We hope to be on-line this weekend.  BSM_NH also has an ongoing problem.  For some reason, exposures shorter than about 2 seconds tend to hang MaximDL with this particular ST-8 camera.  It is intermittent enough that sending the camera back to SBIG for repair is not a good option.  Until we find a consistent failure mode, we will probably restrict exposures to be longer than 2 seconds.

BSM_NM's camera was sent back to SBIG for cleaning, but Bill Stein installed his spare ST-10 for operations starting Feb 13 (thanks, Bill!).  The cleaned original camera was re-installed on March 12.  Earlier (around March 2), cabling was redone to protect the computer, and automatic roof closing was implemented.  During these changes, the focuser has quit moving and so the system is out of focus.  Bill will be going to the observatory soon to diagnose and solve the problem, but until then, BSM_NM is off-line.

TMO61 has been having some mount problems.  If the telescope tracks too far west, it takes too much motor current to return to the meridian.  Jon Holtzman is investigating, but we may limit western data-collection for a while.

All other systems are operational.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484